# Deep Reinforcement Learning: A Convex Optimization Approach

### 29 Feb 2024 v1

@@ 2 Episodic Deep Reinforcement Learning with Convex Optimization
$\min_{\begin{subarray}{c}w\\ \alpha>0\end{subarray}}$ $~{}~{}\left|\sum_{p=1}^{P}D_{p}X(v_{1,p}w_{1,p}-v_{2,p}w_{2,p})-y\right|^{2}$ (5)
@@
$\lambda I\preceq\frac{1}{T}\left(\sum_{t=1}^{T}\begin{bmatrix}X_{t}^{\intercal}\\ -X_{t}^{\intercal}\end{bmatrix}\begin{bmatrix}X_{t}\\ -X_{t}\end{bmatrix}+F\right)$
@@ 4 Proof of Theorem 1
$\min_{\begin{subarray}{c}w\\ \alpha>0\end{subarray}}$ $~{}~{}\left|\sum_{p=1}^{P}D_{p}X(v_{1,p}w_{1,p}-v_{2,p}w_{2,p})-y\right|^{2}$ (7)
@@
$A^{\intercal}A+F$ $=\sum_{t=1}^{T}\sum_{p=1}^{P}[D_{p}]_{tt}\begin{bmatrix}X_{t}^{\intercal}\\ -X_{t}^{\intercal}\end{bmatrix}\begin{bmatrix}X_{t}\\ -X_{t}\end{bmatrix}+F$
$\succeq\sum_{t=1}^{T}\sum_{p=1}^{P}\begin{bmatrix}X_{t}^{\intercal}\\ -X_{t}^{\intercal}\end{bmatrix}\begin{bmatrix}X_{t}\\ -X_{t}\end{bmatrix}+F$
@@ 5 Conclusions
$\|w_{k}-w^{\star}\|\leq C\cdot\frac{\rho}{\lambda T}\|,$

### 07 Mar 2024 v2

@@ 2 Episodic Deep Reinforcement Learning with Convex Optimization
$\min_{\begin{subarray}{c}w\\ v>0\end{subarray}}$ $~{}~{}\left|\sum_{p=1}^{P}D_{p}X(v_{1,p}w_{1,p}-v_{2,p}w_{2,p})-y\right|^{2}$ (5)
@@
$\lambda I\preceq\frac{1}{T}\left(\sum_{t=1}^{T}\begin{bmatrix}X_{t}^{\intercal}\\ -X_{t}^{\intercal}\end{bmatrix}\begin{bmatrix}X_{t}&-X_{t}\end{bmatrix}+F\right)$
@@ 4 Proof of Theorem 1
$\min_{\begin{subarray}{c}w\\ v>0\end{subarray}}$ $~{}~{}\left|\sum_{p=1}^{P}D_{p}X(v_{1,p}w_{1,p}-v_{2,p}w_{2,p})-y\right|^{2}$ (7)
@@
$A^{\intercal}A+F$ $=\sum_{t=1}^{T}\sum_{p=1}^{P}[D_{p}]_{tt}\begin{bmatrix}X_{t}^{\intercal}\\ -X_{t}^{\intercal}\end{bmatrix}\begin{bmatrix}X_{t}&-X_{t}\end{bmatrix}+F$
$\succeq\sum_{t=1}^{T}\sum_{p=1}^{P}\begin{bmatrix}X_{t}^{\intercal}\\ -X_{t}^{\intercal}\end{bmatrix}\begin{bmatrix}X_{t}&-X_{t}\end{bmatrix}+F$
@@ 5 Conclusions
$\|w_{k}-w^{\star}\|\leq C\cdot\frac{\rho}{T},$